# Quantum enhanced ensemble GANs for anomaly detection in continuous biomanufacturing

### v1

@@ 3.1 Experiments with an ensemble of 30 discriminators
Fig.  5: Receiver operating characteristic (ROC) curves of our anomaly detection algorithm based on an ensemble of 30 GAN discriminators. We consider GANs that use a classical Gaussian latent space or a quantum distribution produced by the quantum interference between identical photons. The quantum distribution is generated either by a real or by a simulated quantum processor. The GANs differ only in the process that produced the latent vectors sent to the generator. The solid line and shaded areas correspond respectively to the mean and standard deviation calculated by sampling 30 models 20 times from an ensemble of 120 trained models.
@@
Table 1: Area under the curve (AUC) of the receiver operating characteristic curves of our anomaly detection algorithm for a size-30 bootstrap ensemble.
@@ 3.2 Using a larger ensemble
Fig.  6: ROC curves of our anomaly detection method based on an ensemble of 60 GAN discriminators using different latent spaces.
@@
Table 2: Area under the curve (AUC) of the receiver operating characteristic curves of our anomaly detection algorithm for a size-60 bootstrap ensemble.
@@ 4.1 Advantage from quantum computing
Here, we further discuss the role of quantum computing in helping us achieve improved anomaly detection results. We consider two aspects of this question: whether the quantum computer is required to achieve higher-quality results, and the extent to which the quantum computer could be simulated classically.
First, though the quantum distribution provides a practical and well-studied method to improve the performance of a GAN, it is not the only method available. A wide range of other classical distributions can also be used in a GAN, from Bernoulli distributions 54 to trained distributions using an autoencoder 55. Though bacarreza2025quantum 31 observe an improvement over a wide range of classical distributions, we cannot rule out that some of these methods may yield improved performance on this problem compared to the baseline Gaussian. However, it is impractical and computationally expensive to investigate a wide range of alternative distributions, while the increasing availability of quantum computers means that they are becoming an increasingly practical alternative. As such, in realistic settings where algorithm development time and computation is limited, adopting a quantum distribution as the go-to option for improving performance may be a more effective use of resources than investigating a wider range of classical distributions.
Second, we note that classical methods for simulating photonic quantum processors in the presence of noise have been developed 56, 57. The methods could in principle be used to derive the same improvement in performance that we observe with a real quantum processor using only classical resources. However, these methods are not practical for hybrid quantum/classical machine learning workflows. The mean-field approach of villalonga2021efficient 56 reproduces the low-order statistics of a photonic quantum distribution with a quadratic overhead as opposed to a linear overhead at most with a real system. The approach of oh2024classical 57 was reported to require 9 minutes on 288 GPUs just to set up the simulation. As such, it is unlikely that either approach could run faster than a state-of-the-art photonic quantum processor. Though speed of generating samples is less of a concern when using a static, untrained latent space, this can be a major limitation for more advanced variational settings. In those settings where both the quantum and the classical parts of the algorithm are regularly updated, neither approach thus provides a practical route to simulating these distributions.
@@ 5 Conclusions
In this study, we have introduced a novel framework for unsupervised anomaly detection in continuous biomanufacturing processes, leveraging an ensemble of generative adversarial networks (GANs) enhanced by hybrid quantum-classical approaches. By developing a benchmark dataset using the KTB-1 dynamic simulation model for lovastatin production, we simulated both normal operations and anomalies arising from sudden feedstock variability—a common challenge in bioprocesses. Our ensemble GAN method, trained exclusively on normal data, demonstrated robust early detection capabilities, with receiver operating characteristic (ROC) curves indicating effective discrimination between normal and anomalous regimes. A key innovation lies in the integration of quantum-generated latent distributions, sourced from both simulated and real photonic quantum processors (ORCA Computing PT-2). Comparative experiments revealed that these quantum distributions outperformed classical Gaussian baselines, yielding higher areas under the curve (AUC) values—up to 0.429 for simulated quantum ensembles of 60 discriminators. This improvement underscores the potential of quantum mechanics to enhance generative diversity, mitigating issues like mode collapse and enabling discriminators to learn more informative decision boundaries. Notably, even imperfect real quantum hardware provided advantages over classical methods, validating the practicality of hybrid architectures in near-term applications. This work represents the first application of hybrid quantum-classical GANs for anomaly detection in continuous biomanufacturing, addressing critical needs for robust process monitoring amid nonlinear dynamics and regulatory demands. The framework’s unsupervised nature and scalability with ensemble size position it as a promising tool for integration with model predictive control (MPC) systems, facilitating proactive interventions to minimize yield losses and operational disruptions. While our results highlight quantum advantages, limitations include reliance on small-scale quantum processors and the computational cost of larger ensembles. Future research could explore trained quantum circuits, scalability to fault-tolerant quantum hardware, and real-world validation on industrial datasets. Overall, this study paves the way for quantum-enhanced machine learning to drive advancements in sustainable biomanufacturing, offering a pathway toward more resilient and efficient processes.

### v2

@@ 3.1 Experiments with an ensemble of 30 discriminators
Fig. 5: Receiver operating characteristic (ROC) curves of our anomaly detection algorithm based on an ensemble of 30 GAN discriminators. We consider GANs that use a classical Gaussian latent space or a quantum distribution produced by the quantum interference between identical photons. The quantum distribution is generated either by a real or by a simulated quantum processor. The GANs differ only in the process that produced the latent vectors sent to the generator. The solid line and shaded areas correspond respectively to the mean and standard deviation calculated by sampling 30 models 20 times from an ensemble of 120 trained models.
@@
Table 1: Area under the curve (AUC) of the ROC curve in figure 5, with a size-30 bootstrap ensemble. Perfect detection = 0.5
@@
Random guessing 0.125
@@ 3.2 Using a larger ensemble
Fig. 6: ROC curves of our anomaly detection method based on an ensemble of 60 GAN discriminators using different latent spaces.
@@
Table 2: Area under the curve (AUC) of the ROC curve in figure 6, with a size-60 bootstrap ensemble. Perfect detection = 0.5
@@
Random guessing 0.125
@@ 4 Discussion
4.1 Strengths and weaknesses of a GAN-based approach
Though our GAN-based approach is advantageous in terms of applicability to complex datasets that may arise in a biomanufacturing setting, some disadvantages must be acknowledged. First, using an ensemble of GANs can be computationally intensive, with size-30 and size-60 ensembles considered in our work. Though each model we consider is small and can be trained in under 5 minutes on an Apple M2 chip, scaling to larger models and datasets may be onerous. However, we note that after training, inference can often be performed quickly and cheaply.
We also note that other deep learning based approaches have been developed for anomaly detection and could be applied here. Though GANs for time series anomaly detection such as BeatGAN 59 have been successful, other approaches have been developed such as Anomaly Transformer 56. However, our objective with this work was not to beat existing benchmarks for anomaly detection in biomanufacturing, but rather to develop and investigate a new approach for doing so. Further refining the method and comparing it to other methods on a broader range of problems would be a natural extension of this work.
@@ 4.2 Advantage from quantum computing
Here, we further discuss the role of quantum computing in helping us achieve improved anomaly detection results. Though this current work used a distribution that is native to quantum processors, the small-scale nature of the distribution with 8 photons in 16 qumodes makes it classically simulable. However, scaling this approach to larger-scale datasets is likely to require larger processors that are no longer simulable. For these larger-scale quantum processors, we consider two questions: whether the quantum computer is required to achieve higher-quality results, and the extent to which the quantum computer could be, if not exactly simulated, at least approximated classically.
First, though the quantum distribution provides a practical and well-studied method to improve the performance of a GAN, it is not the only method available. A wide range of other classical distributions can also be used in a GAN, from Bernoulli distributions 10 to trained distributions using an autoencoder 24. Though Bacarreza et al. 6 observe an improvement over a wide range of classical distributions, we cannot rule out that some of these methods may yield improved performance on this problem compared to the baseline Gaussian. However, it is impractical and computationally expensive to investigate a wide range of alternative distributions, while the increasing availability of quantum computers means that they are becoming an increasingly practical alternative. As such, in realistic settings where algorithm development time and computation is limited, adopting a quantum distribution as the go-to option for improving performance may be a more effective use of resources than investigating a wider range of classical distributions.
Second, we note that classical methods for simulating photonic quantum processors in the presence of noise have been developed 52, 39. The methods could in principle be used to derive the same improvement in performance that we observe with a real quantum processor using only classical resources. However, these methods are not practical for hybrid quantum/classical machine learning workflows. The mean-field approach of Villalonga et al. 52 reproduces the low-order statistics of a photonic quantum distribution with a quadratic overhead as opposed to a linear overhead at most with a real system. The approach of Oh et al. 39 was reported to require 9 minutes on 288 GPUs just to set up the simulation. As such, it is unlikely that either approach could run faster than a state-of-the-art photonic quantum processor. Though speed of generating samples is less of a concern when using a static, untrained latent space, this can be a major limitation for more advanced variational settings. In those settings where both the quantum and the classical parts of the algorithm are regularly updated, neither approach thus provides a practical route to simulating these distributions.
@@ 5 Conclusions
In this study, we have introduced a novel framework for unsupervised anomaly detection in continuous biomanufacturing processes, leveraging an ensemble of generative adversarial networks (GANs) enhanced by hybrid quantum-classical approaches. By developing a benchmark dataset using the KTB-1 dynamic simulation model for lovastatin production, we simulated both normal operations and anomalies arising from sudden feedstock variability—a common challenge in bioprocesses. Our ensemble GAN method, trained exclusively on normal data, demonstrated robust early detection capabilities, with receiver operating characteristic (ROC) curves indicating effective discrimination between normal and anomalous regimes.
A key innovation lies in the integration of quantum-generated latent distributions, sourced from both simulated and real photonic quantum processors (ORCA Computing PT-2). Comparative experiments revealed that these quantum distributions outperformed classical Gaussian baselines, yielding higher areas under the curve (AUC) values—up to 0.429 for simulated quantum ensembles of 60 discriminators. This improvement underscores the potential of quantum mechanics to enhance generative diversity, mitigating issues like mode collapse and enabling discriminators to learn more informative decision boundaries. Notably, even imperfect real quantum hardware provided advantages over classical methods, validating the practicality of hybrid architectures in near-term applications.
This work represents the first application of hybrid quantum-classical GANs for anomaly detection in continuous biomanufacturing, addressing critical needs for robust process monitoring amid nonlinear dynamics and regulatory demands. The framework’s unsupervised nature and scalability with ensemble size position it as a promising tool for integration with model predictive control (MPC) systems, facilitating proactive interventions to minimize yield losses and operational disruptions.
While our results highlight quantum advantages, limitations include reliance on small-scale quantum processors and the computational cost of larger ensembles. Future research could explore trained quantum circuits, scalability to fault-tolerant quantum hardware, and real-world validation on industrial datasets. Overall, this study paves the way for quantum-enhanced machine learning to drive advancements in sustainable biomanufacturing, offering a pathway toward more resilient and efficient processes.
@@ Appendix A Data pre-processing
Appendix B Individual ROC curves
Fig. 7: ROC curves of 30 different discriminators used on their own to detect anomalies for (top) simulated quantum and (bottom) Gaussian latent distributions.
To provide insight into the behavior of individual constituents of the GAN ensembles used for anomaly detection, figure 7 shows ROC curves for 30 randomly selected discriminators for two different latent distributions. First, we observe that every discriminator significantly underperforms the ensemble with the ROC curve in figure 5. This shows that the ensembling provides a benefit. Second, we see very heterogeneous performance between different discriminators. Since all the GANs successfully trained to convergence, this suggests that different decision boundaries are indeed learned by the discriminators, and only some of which are useful for the type of anomaly investigates in this work. Third, we see more diversity in discriminator performance with the quantum latent distribution especially at lower false positive rates, with 4 discriminators significantly outperforming the others. This suggests that the quantum latent distribution learned a broader set of decision boundaries, leading to improved performance for a subset of the discriminators on this task.
Appendix C ORCA Computing PT-2 system
The photonic quantum processor used in this work is an ORCA Computing PT-2 system. The PT-2 consists of a time-bin interferometer as shown in figure 4, where a single pulsed quantum light source produces a sequence of regularly spaced photons. These photons interfere with each other inside a sequence of two optical delay lines, where the length of each delay line matches the time delay between photons. This allows the photons to interfere with other photons in subsequent pulses, creating a complex superposition state. A single detector at the output can resolve up to 6 photons in each time bin.
The system we use in this work is hosted at the UK National Quantum Computing Centre. Our experiments use 16 time bin bins, with photons injected at each time bin. Due to system losses of roughly 70%, fewer photons are detected than are sent in. To approximate the ideal simulation where 8 photons are detected, we use "post-selection", where we only return outcomes where at least 8 photons were detected. Other sources of imperfection, such as photon distinguishability, beam splitter inaccuracies or detector dark counts were all found to play a minor role compared to system loss.
For our experiments with a simulated quantum processor, we use the simulation algorithm by Clifford and Clifford 12. We consider an input state with single photons alternating with vacuum in 16 channels, for a total of 8 photons. These simulations assume an ideal quantum processor with no loss and perfect photon number resolution. The circuit we use is the one shown in figure 4, with randomly selected beam splitter parameters.
Appendix D Dataset analysis
Here, we provide additional information concerning the dataset used for this task. All data was generated from the same plant-wide dynamic simulator with identical process noise (Gaussian, $\sigma=0.1$). Figure 8 shows typical trajectories of the 5 features we consider in the data for both the normal data used for training and anomalies in which a 0.5% drop in lactose feed concentration occurred after 10 hours. We observe that for each feature, the anomalous data overlaps significantly with the normal data, with divergence only towards the end of the trajectory. Moreover, the gap between anomalous and normal data at the end of the trajectory is relatively small, with overlap in some the features.
To further quantify the difference between normal and anomalous data, in figure 9 we plot histograms of both types of data along the first two PCA components of the normal data. We observe almost complete overlap in the first PCA component, which captures 95% of the variance in the normal data. A difference between normal and anomalous data only becomes apparent when considering the second PCA component. These results illustrate the importance of learning good decision boundaries in the data. A simple model that learns a decision boundary based on the first PCA component will not identify these anomalies. A more refined model that is able to learn more complex decision boundaries is required. Though in this case a linear model based on the second PCA component in the data would have been successful, in the unsupervised anomaly detection scenario this is not information that would have been available a priori. Using a neural network for this task provides the flexibility to autonomously learn linear or non-linear decision boundaries where appropriate.
In addition to our experiments with a 0.5% disturbance, we also initially investigated a dataset with a 1.5% disturbance. This was however found to be too simple, leading us to focus our efforts instead on the 0.5% dataset.
Fig. 8: Trajectories of the selected features. These correspond to some of the most important propagating control variables and measured variables, and are respectively recycle factor, product concentration after the nanofilter, recycle rate, product concentration after reactor, and lactose concentration after reactor. The blue curves are sample trajectories from the normal dataset, and the red curves are sample trajectories from the anomalous dataset.
Fig. 9: Histograms of the first (top) and second (bottom) PCA components of the normal (blue) and anomalous (red) data.